# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v9

@@ 5. Main Result and Proof
Consider the curvature formula (2.5). The mixed contribution of $T=\omega_{h}+\text{(coexact/exact)}$ to $R^{C}$ has two sources. First, the covariant derivative term $\nabla^{LC}T$ produces operators that, when $X,Y$ lie in $V_{1}$ and $Z$ in $V_{2}$ (or vice versa), insert a mixed $3$-form and hence map between $V_{1}$ and $V_{2}$. Second, the quadratic term $Q_{T}$ contains contractions of two copies of $T$ which, for mixed arguments, again produce off‑diagonal operators. Pointwise at $p$, for each simple mixed component $\alpha_{i}\wedge\beta_{i}$ one can choose $X,Y\in V_{1}$ so that $\alpha_{i}(X,Y)\neq 0$ and $Z\in V_{2}$ so that $\beta_{i}(Z)\neq 0$. Substituting into (2.5) yields an endomorphism with a nontrivial $V_{2}\to V_{1}$ component; dually, for each $\tilde{\alpha}_{j}\wedge\tilde{\beta}_{j}$ one chooses $X\in V_{1}$, $Y,Z\in V_{2}$ to obtain a nontrivial $V_{1}\to V_{2}$ component.
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists an open‑dense subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
@@
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}).$ (5.5)
The map is well‑defined and linear; its image consists precisely of the off‑diagonal curvature operators generated by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the (2,1)(2,1)‑component of $\omega_{h}$.
If $\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$, then the corresponding off‑diagonal curvature contribution vanishes:
@@
Using the product rule $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})$ and the fact that $R^{LC}_{\mathrm{off}}=0$, this condition is equivalent to
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0.$ (5.6)
Separating the bidegrees in (5.6) yields the two independent conditions
$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.7)
Because the forms $\beta_{i}$ (resp. $\alpha_{i}$) are linearly independent on an open‑dense set, we obtain
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}=0$ (5.8)
on that set. By the unique‑continuation principle for harmonic forms (Lemma 5.1), these equalities extend globally, hence
$\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.9)
@@
For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Equation (5.9) gives
$A:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad B:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$. Set $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and $V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2})$. Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{\dim U}\}$ whose classes are a basis of the quotient $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$). By Lemma 5.1, at a generic point $p_{1}\in M_{1}$ the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{\dim U}$ are linearly independent. Since $A$ is parallel, $\nabla^{g_{1}}A\equiv 0$; evaluating at such a $p_{1}$ yields
$\sum_{i>s}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0,$
and the linear independence forces $c_{i}=0$ for all $i>s$.
An identical construction applied to $M_{2}$ (choosing a basis of $V_{\parallel}$ and completing it to a basis of $V$ such that the remaining vectors form a basis of the quotient $V/V_{\parallel}$) implies $c_{i}=0$ for all $i>t$.
@@
The opposite inclusion is immediate: any element of $\mathcal{V}_{2,1}$ that belongs to $\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})$ is a finite sum of simple tensors of parallel forms; consequently its covariant derivative vanishes, and hence it lies in $\ker\Psi_{p}$.
Thus we have proved
@@
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)$ (5.11)
and
@@
Hence, on the open‑dense set $\mathcal{U}$ where the harmonic evaluations are independent, the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
@@
linearly independent off‑diagonal curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra, whence
@@
Consider the one-parameter family of connections
$\nabla^{(s)}=\nabla^{LC}+\frac{s}{2}T.$
Then the off-diagonal curvature satisfies $R_{\mathrm{off}}(s)=s(\nabla^{LC}T)_{\mathrm{off}}+s^{2}Q_{\mathrm{off}}$. Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\mathrm{off}}$ has dimension at least $r^{\sharp}$. The key step is to ensure that the rank of $R_{\mathrm{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\mathrm{off}}$.
The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$. On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$. The rank $r^{\sharp}$ is established by the harmonic component $\omega_{h}$ of $T$. The non-harmonic part $T-\omega_{h}=d\alpha+\delta\beta$ contributes to $A_{1}$, but since the bound is a lower bound, this contribution cannot reduce the rank below $r^{\sharp}$ on the dense set $U$ and is therefore absorbed by the overall estimate $\mathrm{rank}(A_{1})\geq r^{\sharp}$. More formally, a systematic cancellation of the off-diagonal contributions by the non-harmonic part $\eta=d\alpha+\delta\beta$ on an open set would require $\nabla\eta=-\nabla\omega_{h}$ locally. By the Unique Continuation Principle, this identity would extend globally, implying $\omega_{h}]=0$, which contradicts our hypothesis of a non-trivial mixed class. This ensures that the topological information carried by $\omega_{h}$ is rigidly encoded in the curvature.
To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in Remark 5.6).
By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), we can select a sequence of real-analytic product metrics $g_{k}=g_{1,k}\oplus g_{2,k}$ converging to $g$ in the $C^{\infty}$ topology and satisfying
$\dim\mathcal{P}_{k}(M_{i},g_{k})\leq\dim\mathcal{P}_{k}(M_{i},g)\qquad(k=1,2).$
This selection is possible because the dimension function $g^{\prime}\mapsto\dim\mathcal{P}_{k}(M_{i},g^{\prime})$ is upper semicontinuous; consequently, the set $\{g^{\prime}\mid\dim\mathcal{P}_{k}(M_{i},g^{\prime})\leq\dim\mathcal{P}_{k}(M_{i},g)\}$ is open in the $C^{\infty}$ topology. Real-analytic metrics are dense in this open set by Greene–Jacobowitz. For such a sequence we have $\dim\mathcal{K}(g_{k})\leq\dim\mathcal{K}(g)$, and therefore
$r^{\sharp}(g_{k})\;:=\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}(g_{k})\;\geq\;r^{\sharp}(g).$
For each fixed $k$, the curvature tensor $R_{\mathrm{off},k}(s)$ of the connection $\nabla^{(s)}=\nabla^{LC}(g_{k})+\frac{s}{2}T$ is real-analytic in $(s,p)$. On the dense open set $U$ furnished by Lemma 5.1 we have $\operatorname{rank}(A_{1,k})=r^{\sharp}(g_{k})$, where $A_{1,k}=\frac{1}{2}(\nabla^{LC}(g_{k})T)_{\mathrm{off}}$. Hence there exists an $r^{\sharp}(g_{k})\times r^{\sharp}(g_{k})$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ with $M_{k}(s)=s^{r^{\sharp}(g_{k})}D_{k}(s)$ and $D_{k}(0)\neq 0$. By the Identity Principle for Analytic Functions, $D_{k}(s)$ does not vanish identically and its zero set is discrete; consequently, for each $k$ there exists an open–dense subset $U_{k}\subseteq U$ such that
$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k})\qquad\forall p\in U_{k}.$
Let $W:=\bigcap_{k=1}^{\infty}U_{k}$. Since $M$ is compact, it carries a complete Riemannian metric; in a complete metric space, a countable intersection of open dense subsets is dense. Thus $W$ is dense in $U$, hence dense in $M$. For every $p\in W$ and every $k$ we have
$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k}).$
The rank of a continuous family of linear maps is lower semicontinuous with respect to uniform convergence [9, Warner]. Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, for every $p\in W$ we obtain
$\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;\liminf_{k\to\infty}\operatorname{rank}R_{\mathrm{off},k}(p)\;\geq\;\liminf_{k\to\infty}r^{\sharp}(g_{k})\;\geq\;r^{\sharp}(g).$
The set $\{p\in M\mid\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)\}$ is open (by lower semicontinuity) and contains the dense subset $W$; therefore it is all of $M$. Thus $\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)$ for every $p\in M$.
Finally, by Definition 3.2 the off‑diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ is the Lie subalgebra generated by the off‑diagonal curvature operators and their covariant derivatives at $p$. In particular, it contains the linear span of the off‑diagonal curvature operators $R_{\mathrm{off}}(p)$. Therefore,
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;r^{\sharp}(g)\qquad\forall p\in M,$
@@
which is the desired bound. The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$. ∎
@@
The invariance of the lower bound $r^{\sharp}$ is a non‑trivial and geometrically significant property. The bound depends only on two pieces of data:
@@
Explicitly,
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (resp. $\mathcal{V}_{1,2}$) denotes the linear span of the harmonic representatives of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$.
Consequently, $r^{\sharp}$ is invariant under any deformation of the product metric $g=g_{1}\oplus g_{2}$ that preserves the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ ($k=1,2$). Such deformations include, for instance, any variation of $g_{i}$ inside the connected component of the moduli space of metrics with a fixed set of parallel forms.
This stability is especially relevant when at least one factor $M_{i}$ possesses special holonomy (e.g., Calabi–Yau manifolds or flat tori), where the spaces $\mathcal{P}^{k}(M_{i})$ are non‑trivial and the metric $g_{i}$ admits large families of deformations that leave $\mathcal{P}^{k}(M_{i})$ unchanged.
Thus, $r^{\sharp}$ quantifies the part of the topological “mixing strength” of $\omega]$ that cannot be “absorbed” by metric deformations respecting the parallel‑form strata; it captures a genuinely topological obstruction to reducible holonomy.
@@
Remark 5.6 (Analyticity as a tool of proof).
The use of real‑analytic metrics in the proof of Theorem 5.2 is purely instrumental. By the density theorem of Greene–Jacobowitz [4] and the upper semicontinuity of the dimension of the space of parallel forms, one can approximate any smooth product metric $g=g_{1}\oplus g_{2}$ in the $C^{\infty}$-topology by a sequence of real‑analytic product metrics $\{g^{(k)}\}$ satisfying $\dim\mathcal{P}_{\ell}(M_{i},g^{(k)})\leq\dim\mathcal{P}_{\ell}(M_{i},g)$ for $\ell=1,2$.
For each analytic metric $g^{(k)}$ the curvature tensor $R_{\mathrm{off},k}(s)$ depends analytically on the deformation parameter $s$ and on the point $p\in M$; consequently the rank of the off‑diagonal curvature is lower‑semicontinuous and can be controlled by the behaviour of analytic minors (Identity Principle).
Passing to the limit $k\to\infty$ via the lower‑semicontinuity of matrix rank with respect to uniform convergence [9] yields the desired bound for the original $C^{\infty}$ metric $g$. Thus the analytic approximation does not impose any restriction on the result—it merely provides a convenient technical tool for extending the estimate from a generic dense subset to the whole manifold.

### v10

@@ 5. Main Result and Proof
Consider the curvature formula (2.5). The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts.
Our aim is to show that the off‑diagonal holonomy algebra contains at least $r^{\sharp}$ linearly independent directions. These directions will be produced by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the harmonic part $\omega_{h}$ alone. We must verify that the remaining contributions—the part of the linear term coming from $\eta$ and the quadratic term $Q_{T}$—cannot interfere with this lower bound.
Suppose that on some non‑empty open set $V\subset M$, the derivative $\nabla^{LC}\eta$ cancelled enough off‑diagonal directions from $\nabla^{LC}\omega_{h}$ to reduce the rank of $(\nabla^{LC}T)_{\mathrm{off}}$ below $r^{\sharp}$. This would impose linear relations among the covariant derivatives of the harmonic components $\alpha_{i}\wedge\beta_{i}$ on $V$. Applying the analytic‑approximation technique employed in Lemma 5.1 (see also Remark 5.6), we may work with real‑analytic metrics and forms. The relation then becomes an equality between real‑analytic sections, so by the unique continuation principle for analytic sections (Aronszajn [1]) it extends to the whole connected component containing $V$, hence globally. Since the tensors $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})$ are pointwise linearly independent on the open set $\mathcal{U}$ provided by Lemma 5.1, evaluating the global relation at any point of $\mathcal{U}$ forces all coefficients to vanish, contradicting the existence of a non‑trivial cancellation. Hence, on a (possibly smaller) non‑empty open set, the linear term $(\nabla^{LC}T)_{\mathrm{off}}$ still exhibits at least $r^{\sharp}$ independent off‑diagonal directions originating from $\omega_{h}$. Regarding the quadratic term $Q_{T}$, observe that the $r^{\sharp}$ independent operators produced by the linear term belong to the off‑diagonal holonomy algebra by the Ambrose–Singer theorem [2]. The quadratic term contributes additional operators to the same algebra. Since the dimension of a vector space generated by a set of vectors is at least the rank of any linearly independent subset, adding further generators cannot reduce the dimension below $r^{\sharp}$. Consequently, the lower bound obtained from the linear term persists for the full curvature $R^{C}$. Thus, to establish $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ it suffices to analyse the linear term generated by the harmonic component $\omega_{h}$.
To organise these contributions, choose the harmonic forms $\{\alpha_{i}\}$ on $M_{1}$ and $\{\beta_{i}\}$ on $M_{2}$ mutually $L^{2}$-orthonormal within their degrees, and likewise for $\{\tilde{\alpha}_{j}\}$, $\{\tilde{\beta}_{j}\}$. By minimality of the tensor‑rank decomposition, the sets $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ are linearly independent as sections; hence there exists a non‑empty open subset of points $p=(p_{1},p_{2})\in M_{1}\times M_{2}$ where their evaluations $\{\alpha_{i}|_{p_{1}}\}\subset\Lambda^{2}T_{p_{1}}^{*}M_{1}$ and $\{\beta_{i}|_{p_{2}}\}\subset T_{p_{2}}^{*}M_{2}$ are linearly independent. For any such $p$ we pick vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfying $\alpha_{i}(X_{i},Y_{i})|_{p}=1$ and $\beta_{i}(Z_{i})|_{p}=1$.
@@
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}),$ (5.4)
where the vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$ and $Z_{i}\in T_{p_{2}}M_{2}$ satisfy $\alpha_{i}(X_{i},Y_{i})=1$ and $\beta_{i}(Z_{i})=1$.
Claim 5.2.1 (Independence of choices). The subspace $\ker\Psi_{p}\subset\mathcal{V}_{2,1}$ is independent of the choice of normalizing vectors $\{X_{i},Y_{i},Z_{i}\}$ in  5.4.
Proof of Claim.
Writing $\omega_{h}^{2,1}=\sum_{i=1}^{r_{2,1}}\alpha_{i}\wedge\beta_{i}$, the product rule gives
$(\nabla^{LC}\omega_{h})^{\mathrm{off}}=\sum_{i=1}^{r_{2,1}}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i}).$
For vectors $X_{i},Y_{i}\in T_{p_{1}}M_{1}$, $Z_{i}\in T_{p_{2}}M_{2}$ with $\alpha_{i}(X_{i},Y_{i})=1$, $\beta_{i}(Z_{i})=1$, the contraction $\iota_{Z_{i}}\iota_{Y_{i}}\iota_{X_{i}}$ extracts precisely the $i$-th component:
$(\nabla^{LC}\omega_{h})^{\mathrm{off}}(X_{i},Y_{i};Z_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i}).$
Summing over $i$ with coefficients $c_{i}$ yields that $\Psi_{p}(\xi)=0$ is equivalent to the tensor equation
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0,$ (5.5)
which is manifestly independent of the vectors $\{X_{i},Y_{i},Z_{i}\}$. Hence $\ker\Psi_{p}$ is well-defined. ∎
Now, although the right-hand side of   5.4 depends on the choice of vectors, by Claim 5.2.1 the condition $\Psi_{p}(\xi)=0$ is independent of that choice.
If $\xi\in\ker\Psi_{p}$, then by definition
@@
Separating the bidegrees in   5.5 yields the two independent conditions
$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\quad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.6)
Since the harmonic forms $\beta_{i}$ are linearly independent in $\mathcal{H}^{1}(M_{2})$, the set of points where they are pointwise linearly independent is a non‑empty open subset of $M_{2}$ (its complement is the zero locus of certain smooth determinants). Denote this open set by $\mathcal{U}_{2}\subset M_{2}$.
Now fix any point $y_{0}\in\mathcal{U}_{2}$. For this fixed $y_{0}$, the first equation in (5.6) becomes:
$\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)\Bigr)\wedge\beta_{i}(y_{0})=0\quad\text{for all }x\in M_{1}.$
Since the covectors $\beta_{1}(y_{0}),\dots,\beta_{r_{2,1}}(y_{0})$ are linearly independent, this forces
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(x)=0\quad\text{for all }x\in M_{1}.$
Analogously, because the harmonic forms $\alpha_{i}$ are linearly independent in $\mathcal{H}^{2}(M_{1})$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ where they are pointwise linearly independent.
Fix $x_{0}\in\mathcal{U}_{1}$. The second equation in (5.6) then gives for all $y\in M_{2}$:
$\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}(x_{0})\Bigr)\wedge\nabla^{g_{2}}\beta_{i}(y)=0,$
which implies $\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(y)=0$ for all $y\in M_{2}$.
Therefore,
$\eta_{1}:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.7)
@@
For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Then (5.7) holds, i.e.
$\eta_{1}=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\eta_{2}=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$. Define
$U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1}),\qquad V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2}),$
and set
$s=\dim U_{\parallel},\qquad t=\dim V_{\parallel}.$
Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{r_{2,1}}\}$ whose classes form a basis of $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$).
Now apply Lemma 5.1 to the subspace $U\subset\mathcal{H}^{2}(M_{1})$ with its parallel part $U_{\parallel}$ (which has dimension $s$). The lemma guarantees a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$, the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{r_{2,1}}$ are linearly independent.
Since $\eta_{1}$ is parallel, $\nabla^{g_{1}}\eta_{1}\equiv 0$; evaluating at $p_{1}$ we obtain
$\sum_{i=s+1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0.$
The linear independence at $p_{1}$ forces $c_{i}=0$ for all $i=s+1,\dots,r_{2,1}$.
An identical argument applied to $M_{2}$ yields a point $p_{2}\in\mathcal{U}_{2}$ such that
$\sum_{i=t+1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}(p_{2})=0,$
and the linear independence of $\{\nabla^{g_{2}}\beta_{i}(p_{2})\}_{i=t+1}^{r_{2,1}}$ forces $c_{i}=0$ for all $i=t+1,\dots,r_{2,1}$.
@@
The opposite inclusion has already been proved above. Thus
@@
Recall that $s=\dim(U_{\parallel})$ and $t=\dim(V_{\parallel})$. By Lemma 5.1 applied to the subspace $\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ with parallel part of dimension $s$, there exists a non‑empty open set $\mathcal{U}_{1}\subset M_{1}$ such that for every $p_{1}\in\mathcal{U}_{1}$ the covariant derivatives $\nabla^{g_{1}}\alpha_{i}(p_{1})$ (for $i=s+1,\dots,r_{2,1}$) are linearly independent. Similarly, applying Lemma 5.1 to $\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$ with parallel part of dimension $t$ yields a non‑empty open set $\mathcal{U}_{2}\subset M_{2}$ such that for every $p_{2}\in\mathcal{U}_{2}$ the covariant derivatives $\nabla^{g_{2}}\beta_{i}(p_{2})$ (for $i=t+1,\dots,r_{2,1}$) are linearly independent.
Consequently, the product set $\mathcal{U}:=\mathcal{U}_{1}\times\mathcal{U}_{2}$ is a non‑empty open subset of $M=M_{1}\times M_{2}$. For any point $p=(p_{1},p_{2})\in\mathcal{U}$ the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, i.e.
$\dim\operatorname{im}\Psi_{p}=r_{2,1}-\dim\ker\Psi_{p},\qquad\dim\operatorname{im}\widetilde{\Psi}_{p}=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$
Hence, at every $p\in\mathcal{U}$, the off‑diagonal part of the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
@@
linearly independent curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra at $p$. By Definition 3.2 of the off‑diagonal holonomy algebra $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ as the Lie subalgebra generated by all off‑diagonal curvature components and their covariant derivatives, we therefore have
@@
Now let $p_{0}$ be any point in $\mathcal{U}$. Because the manifold $M$ is connected, the holonomy algebras at different points are isomorphic via parallel transport along curves. In particular, the dimension of the off‑diagonal holonomy algebra is constant on $M$. Hence the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$, which holds at $p_{0}$, must hold at every point $p\in M$ and no density requirement is needed. This completes the proof of the lower bound.
@@
The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$. ∎
@@
The invariance of the lower bound $r^{\sharp}$ is a non‑trivial property that holds under the Assumption 2.2. In those cases, the bound depends only on:
@@
Consequently, $r^{\sharp}$ is invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these strata. This stability is relevant for manifolds with special holonomy where $\mathcal{P}_{k}(M_{i})$ are non‑trivial.
@@
Remark 5.6 (Analytic approximation as a tool of proof).
The proofs of Lemma 5.1 and Theorem 5.2 make essential use of real‑analytic approximations of the smooth metric. This technique does not restrict the generality of the results, because real‑analytic metrics are dense in the $C^{\infty}$ topology (Greene–Jacobowitz [4]) and the relevant quantities – the dimension of the space of parallel forms, the harmonic representatives of fixed cohomology classes, and the covariant derivatives of those harmonic forms – depend continuously on the metric in the appropriate topologies.
In Lemma 5.1 the reduction to an analytic metric allows us to apply the identity principle for analytic functions: if an analytic section of a vector bundle vanishes on an open set, it vanishes identically. This principle is used to show that a relation among the covariant derivatives of harmonic forms on an open set forces the existence of a parallel combination of the forms, contradicting the hypothesis of linear independence modulo parallel forms. For a smooth metric, we approximate by analytic metrics, prove the existence of a non‑empty open set of maximal rank for each approximation, and then pass to the limit using $C^{1}$ convergence of the covariant derivatives.
In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1. Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows first on that open set, and then on the whole manifold because the dimension of the holonomy algebra is constant on connected components (Ambrose–Singer [2]).
Thus analyticity serves solely as a convenient tool to handle the delicate rank conditions in Lemma 5.1; the final statements are valid for arbitrary smooth product metrics and do not require any analyticity assumption on the metric or the torsion.